Harris Company manufactures and sells a single product.


Required:
1.

A partially completed schedule of the company’s total and per unit costs over the relevant range of 55,000 to 95,000 units produced and sold annually is given below: Complete the schedule of the company’s total and unit costs. Round the variable cost and fixed cost to 2 decimal places.)

Units Produced and Sold
55,000 75,000 95,000
Total costs:
Variable costs $181,500
Fixed costs 370,000
Total costs $551,500
Cost per unit:
Variable cost
Fixed cost
Total cost per unit

     

2.

Assume that the company produces and sells 85,000 units during the year at a selling price of $9.23 per unit. Prepare a contribution format income statement for the year. (Do not round intermediate calculations.)

Harris Company
Contribution Format Income Statement
Answer #1
Units Produced and Sold
55,000 75,000 95,000
Total costs:
Variable costs $181,500 75,000 x 3.3 = 247,500 95,000 x 3.3 = 313,500
Fixed costs 370,000 370,000 370,000
Total costs $551,500 $617,500 $683,500
Cost per unit:
Variable cost 181,500/55,000 = 3.3 3.3 3.3
Fixed cost 370,000/55,000 =6.73 370,000/75,000 = 4.93 370,000/95,000 = 3.89
Total cost per unit 10.03 8.23 7.19

Formula used:

Fixed cost per unit = Total fixed cost / Number of units

Variable cost per unit = Total variable cost / Number of units

Total variable cost = Variable cost per unit x Number of units

Harris Company
Contribution Format Income Statement
Sales ( 85,000 x 9.23) 784,550
Variable cost (85,000 x 3.3) 280,500
Contribution margin 504,050
Fixed cost -370,000
Operating Profit $134,050